OpenStudy (anonymous):

How do you integrate (54-9y)^(1/2) ?

OpenStudy (anonymous):

\[\int\limits(59-9y)^{1/2} dx\] Do a u-sub: u=59-9y

OpenStudy (espex):

Which means du will be...?

OpenStudy (anonymous):

I don't remember how to use the chain rule for integration....du= -9...

OpenStudy (anonymous):

You can drag the -9 out to the front because it's just a constant, so you end up with \[(1/-9)\int\limits u^{1/2}du\]

OpenStudy (anonymous):

drag the (-1/9)* out front, not -9